### v1

@@ 6.3 Comparison to previous works
Third, TP17 is a pixel-by-pixel analysis where additional information is introduced in the model by assuming similar $v_{\rm LOS}$ for adjacent pixels. The HOG method uses the gradient orientation, a non-local quantity calculated over the angular scale set by the Gaussian derivative kernel size. Then, it accumulates the statistics over the 10$\times$ 10 region to obtain the relation between $d$ and $v_{\rm LOS}$. Thus, TP17 obtains a higher-resolution reconstruction of PPDV than the HOG method at the expense of having to assume priors for the LOS velocity and spatial and velocity coherence. The higher angular resolution in TP17 favors the comparison with the VLBI observations. Still, it can be biased toward identifying the most massive object along a sightline and associating it with the most massive object for the same position in the line emission data. This would correctly assign a distance and velocity to a dense object like a giant molecular cloud (GMC), but produce a less accurate solution for less concentrated neutral gas.
Figure 20 shows the LOS velocity reconstruction obtained in TP17 for the region within 1250 pc from the Sun. From visual comparison with Fig. 11, it is evident that TP17 also recovers the quadrupolar $v_{\rm LOS}$ pattern resulting in circular motions around the Galactic center. However, the amplitude of the reconstructed $v_{\rm LOS}$ in TP17 is much higher than the values we report in Fig. 11, reaching maxima and minima of up to $\pm$40 km s-1. These large residual velocities are beyond the range of our HOG-based reconstruction, given our input range $-25$$<$$v_{\rm LOS}$$<$,25 km s-1. They are recovered using the broader input $v_{\rm LOS}$ range considered in App. A.3.3, but at the expense of a higher chance correlation that biases the HOG results toward higher $v_{\rm LOS}$.
There are only five VLBI maser observations within the ISM volume considered in this paper. Figure 21 shows the difference between the $v_{\rm LOS}$ reconstructions presented in Fig. 11 and TP17 and the observations for those five sources. For our reconstruction, we considered the difference between the nearest distance channel with an assigned $v_{\rm LOS}$ and the position of the maser. The $v_{\rm LOS}$ differences between the masers and the TP17 reconstruction are positive and negative and characterized by a standard deviation of roughly 6.7 km s-1. The most significant discrepancy is for the source G176.51+00.20, where the difference is around --14 km s-1. The $v_{\rm LOS}$ differences between the masers and our reconstructions are also in either direction. They are characterized by standard deviations around 8.8 and 12 km s-1for Hi and CO, respectively. The most considerable difference is also for source G176.51+00.20, around +16 km s-1.
@@
The HOG-based $v_{\rm LOS}$ reconstruction requires the averaging over large sky areas. Thus, it is expected that its results show less agreement with the masers results than the TP17 reconstruction, which has an angular resolution much closer to the maser observations. Yet, Fig. 21 shows that the discrepancy between the maser and HOG $v_{\rm LOS}$for CO is less than km s-1 for most of the sources, and in the case of G090.21+02.11090.2102.11090.21+02.11090.21 + 02.11 it is closer than TP17. The fact that the HOG $v_{\rm LOS}$for Hi is systematically further from the maser values reflects that it samples more diffuse gas, which is not necessarily anchored to the denser gas producing the maser emission. Given the low number of maser sources to anchor both reconstructions, the comparison with the VLBI results is limited. Thus, we focus on comparing our results and those in TP17. We reserve further comparison between the HOG and the masers for Appendix D.
The differences in the HOG and TP17 $v_{\rm LOS}$ reconstruction methods provide several potential causes for the discrepancies in the results reported in Fig. 11 and Fig. 20. It is plausible that the high $v_{\rm LOS}$ values in TP17 for the range $d$$<$ 1250 kpc are due to features not resolved in the Edenhofer et al. (2024) 3D extinction models or washed away by the spatial averaging required to compute the HOG. Peek et al. (2022) presented a detailed study of a collection of CO clouds in the range 135$<$$l$$<$ 160 using the 3D dust reconstructions from Green et al. (2019) and identified a few objects within $d$$<$ 1250 kpc with $v_{\rm LOS}$$<$$-30$ km s-1; namely the clouds identified with numbers 3397, 3430, 4338, 5935, 6078, and 6142 in the Miville-Deschênes et al. (2017) catalog. The angular size of these clouds ranges between 0 .23 and 0 .67, so it is unlikely that they are the dominant feature of the average across each 5$\times$ 5 distance channels in which the HOG is computed. In that case, the discrepancy is introduced by the different scales sampled by each method.
The HOG-based kinetic tomography employs the information contained in the plane-of-the-sky distribution of the tracers to distinguish the motions in Hi and CO. In that sense, it provides additional information to break the degeneracies in the inversion problem tackled with the TP17. Yet, it is limited by the resolution of the 3D dust reconstruction and the spatial averaging. Thus an optimal approach would be the combination of both methods, using the pixel-by-pixel inversion in TP17 and the HOG-based spatial correlation to regularize the solutions and solve the $v_{\rm LOS}$ reconstruction ambiguities. This merging of techniques is not straightforward and falls beyond the scope of this paper, but should bridge between our results and those in TP17.
6.3.1 Velocities offsets between Hi and CO
@@ 6.4.1 Kinetic energy density
We used our streaming motion estimates to calculate the kinetic energy densities, $E_{\rm k}$, across tiles in the studied volume. These correspond to just one component of the velocity field. However, until new revolutionary techniques or observational methods are invented, this is the only component of the gas velocity we can measure. Assuming equipartition and isotropy, these estimates should be roughly within a factor of three from the three-dimensional values.
@@
The mean $E_{\rm k}$ values of 0.68 and 0.18 eV/cm3 obtained for Hi and CO are comparable to the 0.22 eV/cm3 obtained by assuming standard conditions for the Solar neighborhood (Draine, 2011). Assuming equipartition between the three components of the velocity vector, these estimates should be roughly within a factor of three from the total values. It is a known fact that previous generic estimates of the kinetic energy density are close to the thermal energy density, estimated from the distribution of thermal pressure throughout the ISM, 0.49 eV/cm3 (Jenkins & Tripp, 2011). They are also within a factor of a few from the cosmic ray energy density, 1.39 eV/cm3 (Webber & Yushak, 1983), and the magnetic energy density derived from the Hi Zeeman effect observations, 0.89 eV/cm3 (Heiles & Crutcher, 2005). Finally, they are also close to the energy densities in starlight and far-infrared radiation from dust, 0.54 and 0.31 (see, chapter 12 in Draine, 2011). Our analysis provides a new and independent estimate of the kinetic energy density, demonstrating that this near equipartition is also found in the local ISM.
@@
Still, a novelty introduced by our analysis is the mapping of the distribution of this quantity, which shows maxima a factor of ten and more above the mean value for particular regions, as illustrated in Fig. 15. The distribution of $E_{\rm k}$ across heliocentric distances, presented in the top panel of Fig. 16, shows overdensities concentrated at lower distances. This is most likely produced by the sampling of different angular scales across distances. That is, nearby distance channels have higher spatial resolutions, which resolve dense structures on spatial scales that are not sampled in the most distance channels. This results in a concentration of higher densities at a lower heliocentric radius that is reflected in our $E_{\rm k}$ estimates.
@@
The distance bias is marginalized in the $E_{\rm k}$ azimuthal profiles, presented in the bottom panel of Fig. 16. There it is evident that $E_{\rm k}$ overdensities of a factor of a few are found toward 40$<$$l$$<$ 90 in Hi and CO, around $l$$\approx$ 170 and 270 in CO, and 320$<$$l$$<$ 340 in Hi. The first $l$ range coincides with the position of the Radcliffe wave and the Split, as illustrated in Fig. 3. The $l$$\approx$ 270 feature corresponds to the location of the Vela C MC, also shown in Fig. 3, that is representative of the Vela GMC complexes, which host intermediate-mass star formation, up to early B-type, late O-type stars (Murphy & May, 1991; Netterfield et al., 2009; Massi et al., 2019; Hottier et al., 2021). The region around 320$<$$l$$<$ 340 covers the position of the Ara OB1 association ($l$$\approx$ 336 .7; $b$$\approx$--1 .6), a stellar group extended across 1 on the plane of the sky, where the velocity distribution of Hi has been suggested to correspond to an expanding structure (Arnal et al., 1987). It is unlikely that any of these individual regions by themselves would produce the energy overdensities we are finding in Hi. However, they may be guideposts to larger-scale energy input into the ISM.
@@
Mechanical and radiative energy input from high-mass stars is one of the main sources of kinetic energy and ISM turbulence (see, for example, Hennebelle & Falgarone, 2012; Klessen & Glover, 2016). The largest contribution from high-mass stars to interstellar turbulence most likely comes from SN explosions. Numerical simulations of a single supernova remnant (SNR) expanding in an inhomogeneous density field for different ambient medium’s densities, metallicities, and structures indicate thermal energy inputs of up to $10^{51}$ erg (Martizzi et al., 2015; Iffrig & Hennebelle, 2015). The efficiency of energy transfer from supernova blast waves to the interstellar gas depends on many factors, but it has been estimated to be around 10% (Norman & Ferrara, 1996; Thornton et al., 1998).
@@
The total amount of energy in the regions with energy densities in Hi above the mean is around 8.5 $\times$$10^{51}$ erg. Given the energy input and energy transfer efficiency described above, it would take roughly 85 SNe to introduce that energy into the studied volume. Recently, Swiggum et al. (2024) used Gaia astrometry and other large-scale spectroscopic surveys to show that around 57% of the young cluster within one kiloparsec from the Sun arise from three distinct spatial volumes and these families produced more than 200 SNe in the past 30 Myr, including the sources of the Local Bubble and GSH 238+00+09. This number of SNe is roughly consistent with the 15 Myr-1 SNe for gas at surface densities compared to the local ISM, around 10 M/pc2 (Walch et al., 2015). Given the 12.1 km s-1 velocity dispersion, turbulent mixing would have spread this input energy covering scales of roughly 400 pc in 30 Myr. Thus, SN feedback may be a plausible candidate for the source of the energy overdensities reported in Fig. 15, but it does not entirely explain its distribution.
@@
Figure 24: Distribution of stars in the Zari et al. (2021) catalog limited to the range $b$$<$ presented in Cartesian coordinates (left) and projected into the polar grid introduced in Fig. 3.
In the absence of a direct tracer to locate the SN energy input in the local ISM, our best proxy for the potential location of SNe is the distribution of the current population of high-mass stars. One of the state-of-the-art catalogs of OBA-type stars in the Solar vicinity is presented in Zari et al. (2021). Although this catalog is not complete, due, for example, to the saturation of Gaia for bright sources, it provides a global view of the potential stellar energy input into the ISM. Figure 24 shows the distribution of OBA stars projected into the same grid used for the streaming motion reconstructions. It is evident that the energy overdensities found toward 320$<$$l$$<$ 340 in Fig.15 are in the vicinity of an OBA-type stellar overdensity, making plausible the argument of SNe or other types of stellar feedback as the source of the kinetic energy accumulation. However, this argument does not hold in the region toward 40$<$$l$$<$ 90.
@@
A second scenario is that the extinction within the Split and the Radcliffe Wave hides the stellar population producing the SN feedback. Redding by dust extinction complicates the identification of OBA-type stars, thus these catalog are incomplete toward high-extinction regions. Zari et al. (2021) identified in their figure 14 that the dust density in the range 40$<$$l$$<$ 90 is higher than the stellar density, a fact that they interpreted as the result of the decoupling of gas and stars after the passage of the spiral arm. Testing whether this is the case is problematic because it would involve measuring the plane-of-the-sky gas motions with the same accuracy as we do for stellar motions, which is beyond the reach of current observational techniques. Yet, it is unlikely that the dust in the Split and the Radcliffe Wave are hiding a prominent overdensity of high-mass stars such as that shown in Fig. 24 for distances between 900 and 1250 kpc toward the fourth Galactic quadrant.
@@ 6.4.2 Momentum density
We found standard deviations of the LOS momentum around 0.28 and 0.06 $\times$$10^{43}$ dyn s. These are roughly comparable to the radial momentum input from a single SNR, between approximately 3 $\times$$10^{42}$ and 3 $\times$$10^{43}$ dyn s for forward shock positions between 2 and 12 pc from the explosion (Martizzi et al., 2015). This does not imply that all of the peak $p$ estimated in this paper are produced by SNe but provides a plausible explanation for the diverging motions observed toward the North America MC, which happens along the axis of the Radcliffe wave.
Matter accumulation along the Radcliffe Wave will likely produce star formation that injects energy and momentum into the ISM. However, the accumulation of matter that formed the Radcliffe Wave may be the product of the large-scale Galactic dynamics. Thus, we are again faced with the dichotomy between an SN-dominated and large-scale-forcing-dominated scenario, as we were in the case of the energy distribution over the whole studied volume. Without further evidence, we must conclude that our reconstruction of the momentum density distribution is most likely a combination of both effects.

### v2

@@ 6.3 Comparison to previous works
Third, TP17 is a pixel-by-pixel analysis where additional information is introduced in the model by assuming similar $v_{\rm LOS}$ for adjacent pixels. The HOG method uses the gradient orientation, a non-local quantity calculated over the angular scale set by the Gaussian derivative kernel size. Then, it accumulates the statistics over the 10$\times$ 10 region to obtain the relation between $d$ and $v_{\rm LOS}$. Thus, TP17 obtains a higher-resolution reconstruction of PPDV than the HOG method at the expense of having to assume priors for the LOS velocity and spatial and velocity coherence. The higher angular resolution in TP17 favors the comparison with the VLBI observations. Still, it can be biased toward identifying the most massive object along a LOS and associating it with the most massive object for the same position in the line emission data. This would correctly assign a distance and velocity to a dense object like a giant molecular cloud (GMC), but produce a less accurate solution for less-concentrated neutral gas, which is more exposed to the acceleration by SN feedback and deviations from pure Galactic rotation (Iffrig & Hennebelle, 2015; Walch et al., 2015).
Figure 20 presents the LOS velocity reconstruction obtained in TP17 for the region within 1250 pc from the Sun. From visual comparison with Fig. 11, it is evident that TP17 also recovers the quadrupolar $v_{\rm LOS}$ pattern resulting in circular motions around the Galactic center. However, the amplitude of the reconstructed $v_{\rm LOS}$ in TP17 is much higher than the values we report in Fig. 11, reaching maxima and minima of up to $\pm$40 km s-1. These large residual velocities are beyond the range of our HOG-based reconstruction, given our input range $-25$$<$$v_{\rm LOS}$$<$,25 km s-1. They are recovered using the broader input $v_{\rm LOS}$ range considered in App. A.3.3, but at the expense of a higher chance correlation that biases the HOG results toward higher $v_{\rm LOS}$.
There are only five VLBI maser observations within the ISM volume considered in this paper. Figure 21 shows the difference between the $v_{\rm LOS}$ reconstructions presented in Fig. 11 and TP17 and the observations for those five sources. For our reconstruction, we considered the difference between the nearest distance channel with an assigned $v_{\rm LOS}$ and the position of the maser. The $v_{\rm LOS}$ differences between the masers and the TP17 reconstruction are positive and negative and characterized by a standard deviation of roughly 6.7 km s-1. The most significant discrepancy is for the source G176.51+00.20, where the difference is around --14 km s-1. The $v_{\rm LOS}$ differences between the masers and our reconstructions are also in either direction. They are characterized by standard deviations around 13.5 and 11.3 km s-1for Hi and CO, respectively. The most considerable difference is for source G14.33-00.64, where the lack of morphological correlation between CO and 3D dust limits the $v_{\rm LOS}$ reconstruction.
@@
The HOG-based $v_{\rm LOS}$ reconstruction requires the averaging over large sky areas. Thus, its results are expected to show less agreement with the masers results than the TP17 reconstruction, which has an angular resolution much closer to the VLBI observations. Yet, Fig. 21 shows that the discrepancy between the maser and HOG $v_{\rm LOS}$ for CO is less than a few km s-1 for most of the sources, and in the case of G090.21+02.11 it is closer than TP17. The fact that the HOG $v_{\rm LOS}$ for Hi is systematically further from the maser values reflects that it samples more diffuse gas, which is not necessarily anchored to the denser gas producing the maser emission. Given the low number of maser sources to anchor both reconstructions, the comparison with the VLBI results is limited. Thus, we compare our results to those in TP17. We reserve further comparison between the HOG and the masers for Appendix E.
The differences in the HOG and TP17 $v_{\rm LOS}$ reconstruction methods provide several potential causes for the discrepancies in the results reported in Fig. 11 and Fig. 20. It is plausible that the high $v_{\rm LOS}$ values in TP17 for the range $d$$<$ 1250 kpc are due to features not resolved in the Edenhofer et al. (2024) 3D extinction models or washed away by the spatial averaging required to compute the HOG. Peek et al. (2022) presented a detailed study of a collection of CO clouds in the range 135$<$$l$$<$ 160 using the 3D dust reconstructions from Green et al. (2019) and identified a few objects within $d$$<$ 1250 kpc with $v_{\rm LOS}$$<$$-30$ km s-1; namely the clouds identified with numbers 3397, 3430, 4338, 5935, 6078, and 6142 in the Miville-Deschênes et al. (2017) catalog. The angular size of these clouds ranges between 0 .23 and 0 .67, so it is unlikely that they are the dominant feature of the average across each 10$\times$ 10 distance channels in which the HOG is computed. In that case, the discrepancy is introduced by the different scales sampled by each method.
The HOG-based kinetic tomography employs the information contained in the plane-of-the-sky distribution of the tracers to distinguish the motions in Hi and CO. It provides additional information to break the degeneracies in the inversion problem tackled with the TP17. Yet, it is limited by the resolution of the 3D dust reconstruction and the spatial averaging. Thus, an optimal approach is the combination of both methods, using the pixel-by-pixel inversion in TP17 and the HOG-based spatial correlation to regularize the solutions and solve the LOS reconstruction ambiguities. This merging of techniques is not straightforward and falls beyond the scope of this paper, but should bridge between our results and those in TP17.
@@ 6.4.1 Kinetic energy density
We used our streaming motion estimates to calculate the kinetic energy densities, $E_{\rm k}$, across tiles in the studied volume. These correspond to just one component of the velocity field. However, until new revolutionary techniques or observational methods are invented, this is the only component of the gas velocity we can measure.
The mean $E_{\rm k}$ values of 0.11 and 0.04 eV/cm3 obtained for Hi and CO are comparable to the 0.22 eV/cm3 obtained by assuming standard conditions for the Solar neighborhood (Draine, 2010). Assuming equipartition between the three components of the velocity vector, these estimates should be roughly within a factor of three from the total values.
@@
It is a known fact that previous generic estimates of the kinetic energy density are close to the thermal energy density, estimated from the distribution of thermal pressure throughout the ISM, 0.49 eV/cm3 (Jenkins & Tripp, 2011). They are also within a factor of a few from the cosmic ray energy density, 1.39 eV/cm3 (Webber & Yushak, 1983), and the magnetic energy density derived from the Hi Zeeman effect observations, 0.89 eV/cm3 (Heiles & Crutcher, 2005). Finally, they are also close to the energy densities in starlight and far-infrared radiation from dust, 0.54 and 0.31 (see, chapter 12 in Draine, 2010). Our analysis provides a new and independent estimate of the kinetic energy density, demonstrating that this near equipartition is also found in the local ISM.
@@
Still, a novelty introduced by our analysis is the mapping of the distribution of this quantity, which shows maxima a factor of ten and more above the mean value for particular regions, as illustrated in Fig. 15. The distribution of $E_{\rm k}$ across heliocentric distances, presented in the top panel of Fig. 16, displays overdensities concentrated at lower distances. This is most likely produced by the sampling of different angular scales across distances. That is, nearby distance channels have higher spatial resolutions, which resolve dense structures on spatial scales that are not sampled in the most distance channels. This results in a concentration of higher densities at a lower heliocentric radius that is reflected in our $E_{\rm k}$ estimates.
The potential distance bias is marginalized in the $E_{\rm k}$ azimuthal profiles, presented in the bottom panel of Fig. 16. There it is evident that $E_{\rm k}$ overdensities of a factor of a few are found toward 40$<$$l$$<$ 90 in Hi and CO, around $l$$\approx$ 170 and 270 in CO, and 320$<$$l$$<$ 340 in Hi. The first $l$ range coincides with the position of the Radcliffe wave and the Split, as illustrated in Fig. 3.
The Radcliffe wave comprises the majority of nearby star-forming regions and contains about three million solar masses of gas (Alves et al., 2020). Its existence was not among the premises of our study, but it is easily distinguishable in the Hi $E_{\rm k}$ distribution presented in Fig. 15. Konietzka et al. (2024) used the motion of young stellar clusters to identify oscillatory motions with an amplitude of 14 km s-1 along the Radcliffe wave. The energy overdensities we identified toward this structure are likely related to the oscillatory motions. Konietzka et al. (2024) argued that “a superposition of feedback-driven structures could reproduce the observed wavelength and amplitude of the wave”. The expanding motions in Hi identified in Fig. 23 potentially be remnants of such an input from stellar feedback, but further comparison calls for a study of the Radcliffe wave in it full extent and not just in the region within $|b|$$\leq$ 5.
@@
The $l$$\approx$ 270 energy-overdensity corresponds to the location of the Vela C MC, also shown in Fig. 3, that is representative of the Vela GMC complexes, which host intermediate-mass star formation, up to early B-type, late O-type stars (Murphy & May, 1991; Netterfield et al., 2009; Massi et al., 2019; Hottier et al., 2021). The region around 320$<$$l$$<$ 340 covers the position of the Ara OB1 association ($l$$\approx$ 336 .7; $b$$\approx$--1 .6), a stellar group extended across 1 on the plane of the sky, where the velocity distribution of Hi has been suggested to correspond to an expanding structure (Arnal et al., 1987). It is unlikely that these individual star-forming regions would produce the energy overdensities we find in Hi. However, they may be guideposts to larger-scale energy input into the ISM.
@@
Mechanical and radiative energy input from high-mass stars is one of the primary sources of kinetic energy and ISM turbulence (see, for example, Hennebelle & Falgarone, 2012; Klessen & Glover, 2016). The largest contribution from high-mass stars to interstellar turbulence most likely comes from SN explosions, although recent studies indicate that the mechanical energy injected by the supernovae alone is not sufficient to explain the kinetic energy of the ionized gas in nearby galaxies (Egorov et al., 2023). Numerical simulations of a single supernova remnant (SNR) expanding in an inhomogeneous density field for different ambient medium’s densities, metallicities, and structures indicate thermal energy inputs of up to $10^{51}$ erg (Martizzi et al., 2015; Iffrig & Hennebelle, 2015). The efficiency of energy transfer from supernova blast waves to the interstellar gas depends on many factors, but it has been estimated to be around 10% (Norman & Ferrara, 1996; Thornton et al., 1998).
@@
The total amount of kinetic energy in the studied volume is around 1.7 $\times$$10^{52}$ erg. The total amount of kinetic energy excess relative to a background level, calculated by subtracting the mean from each reconstruction element, is 1.2 $\times$$10^{52}$ erg. Given the energy input and energy transfer efficiency described above, it would take roughly 120 SNe to introduce that energy into that volume. Recently, Swiggum et al. (2024) used Gaia astrometry and other large-scale spectroscopic surveys to show that around 57% of the young stellar clusters within one kiloparsec from the Sun arise from three distinct spatial volumes and these families produced more than 200 SNe in the past 30 Myr, including the sources of the Local Bubble and GSH 238+00+09. This number of SNe is roughly consistent with the 15 Myr-1 SNe for gas at surface densities compared to the local ISM, around 10 M/pc2 (Walch et al., 2015). Given the 10.8 km s-1 velocity dispersion, turbulent mixing would have spread this input energy covering scales of roughly 400 pc in 36 Myr. Thus, SN feedback is a plausible candidate for the source of the energy overdensities reported in Fig. 15, but it does not entirely explain its distribution.
@@
Figure 24: Distribution of O-, B-, and A-type stars in the Zari et al. (2021) catalog limited to the range $b$$<$ 5 presented in Cartesian coordinates (left) and projected into the polar grid introduced in Fig. 3 (right).
In the absence of a direct tracer to locate the SN energy input in the local ISM, our best proxy for the potential location of SNe is the distribution of the current population of high-mass stars. One of the state-of-the-art catalogs of O-, B-, and A-type (OBA) stars in the Solar vicinity is presented in Zari et al. (2021). Although this catalog is not complete, due, for example, to the saturation of Gaia for bright sources, it provides a global view of the potential stellar energy input into the ISM. Figure 24 shows the distribution of OBA stars projected into the same grid used for the streaming motion reconstructions. It is evident that the energy overdensities found toward 320$<$$l$$<$ 340 in Fig.15 are in the vicinity of an OBA-type stellar overdensity, making plausible the argument of SNe or other types of stellar feedback as the source of the kinetic energy accumulation. However, this argument does not hold in the region toward 40$<$$l$$<$ 90.
@@
A second scenario is that the extinction within the Split and the Radcliffe Wave hides the stellar population from which the SN feedback originated. Redding by dust extinction complicates the identification of OBA stars, making catalogs incomplete toward high-extinction regions. Zari et al. (2021) identified in their figure 14 that the dust density in the range 40$<$$l$$<$ 90 is higher than the stellar density, a fact that they interpreted as the result of the decoupling of gas and stars after the passage of the spiral arm. Testing whether this is the case is problematic because it would involve measuring the plane-of-the-sky gas motions with the same accuracy as we do for stellar motions, which is beyond the reach of current observational techniques. Yet, it is unlikely that the dust in the Split and the Radcliffe Wave are hiding a prominent overdensity of high-mass stars such as that shown in Fig. 24 for distances between 900 and 1250 kpc toward the fourth Galactic quadrant.
@@ 6.4.2 Momentum density
Matter accumulation along the Radcliffe Wave may produce star formation that injects energy and momentum into the ISM. However, the initial accumulation of matter that formed the Radcliffe Wave is probably the product of the large-scale Galactic dynamics. Thus, we are again faced with the dichotomy between an SN-dominated and large-scale-forcing-dominated scenario, as we were in the case of the energy distribution over the whole studied volume. Without further evidence, we must conclude that our reconstruction of the momentum density distribution is most likely a combination of both effects.